# ETGL-DDPG: A Deep Deterministic Policy Gradient Algorithm for Sparse Reward Continuous Control

### v2

@@ 1 Introduction
Deep deterministic policy gradient (DDPG) (Lillicrap et al., 2015) is one of the representative algorithms for reinforcement learning (RL) (Sutton & Barto, 2018), alongside other prominent approaches (Haarnoja et al., 2018; Fujimoto et al., 2018; Andrychowicz et al., 2017). In recent years, DDPG has served as the backbone algorithm for introducing novel ideas in robotics and RL Barth-Maron et al. (2018); Pan et al. (2020); Liu et al. (2023); Wang et al. (2023); Tiapkin et al. (2024). While DDPG demonstrates strong performance in continuous control tasks with dense reward signals (Duan et al., 2016; Kiran et al., 2021), its effectiveness diminishes significantly in sparse-reward settings where rewards are only observed upon reaching the goal (Matheron et al., 2019; Luck et al., 2019).
@@
In this paper, we enhance DDPG (Lillicrap et al., 2015) to address all three aforementioned problems. Our first contribution is $\epsilon t$-greedy, a new temporally version of $\epsilon$-greedy that utilizes a light-weight search procedure, similar to Laud (2004), to enable more directional exploration based on the agent’s previous experience data. We show that similar to $\epsilon z$-greedy, $\epsilon t$-greedy has polynomial sample complexity in related parameters of the MDP. Our second contribution is a new goal-conditioned dual replay buffer (GDRB), that uses two replay buffers along with an adaptive sampling strategy to differentiate goal-reached and goal-not-reached experience data. These two buffers differ in retention policy, size, and the transitions they store. Our third enhancement is to replace the one-step update in DDPG with the longest $n$-step return for all transitions in an episode. Figure 1 illustrates the innovations of ETGL-DDPG. In Section 4, we evaluate the performance of ETGL-DDPG through extensive experiments on 2D and 3D continuous control benchmarks. As the proposed strategies are orthogonal, we show that each of the three strategies individually improves the performance of DDPG via a thorough ablation study. Furthermore, ETGL-DDPG outperforms current state-of-the-art methods across all tested environments.
@@ 2.1 Deep Deterministic Policy Gradient (DDPG)
To ease presentation, we adopt our notation with explicit reference to the goal state for both the critic and the actor networks in DDPG. DDPG maintains an actor $\mu(s,s_{g})$ and a critic $Q(s,a,s_{g})$. The agent explores the environment through a stochastic policy $a\sim\mu(s,s_{g})+w$, where $w$ is a noise sampled from a normal distribution or an Ornstein-Uhlenbeck process (Uhlenbeck & Ornstein, 1930). To update both actor and critic, transition tuples are sampled from a replay buffer to perform a mini-batch gradient descent. The critic is updated by a loss $L$,
@@
where $y_{t}=r_{t}+\gamma Q^{\prime}(s_{t+1},\mu^{\prime}(s_{t+1},s_{g}),s_{g})$. $Q^{\prime}$ and $\mu^{\prime}$ are the target critic and actor, respectively; their weights are soft-updated to the current weights of the main critic and actor, respectively. The actor is updated by the deterministic policy gradient algorithm (Silver et al., 2014) to maximize the estimated Q-values of the critic using loss $-\mathbb{E}_{s}[Q(s,\mu(s,s_{g}),s_{g})]$.
@@ 3.1 $\epsilon t$-Greedy: Exploration With Search
To justify this exploration method, we adopt the conditions outlined in Liu & Brunskill (2018) to validate the sample efficiency of $\epsilon t$-greedy. We begin by introducing the relevant terms and then present the main theorem. Detailed definitions and proofs are provided in Appendix A.1. The key idea is to define a measure that captures the concept of visiting all state-action pairs, as outlined in Definition 1.
@@
The covering length (Even-Dar & Mansour, 2004) represents the minimum number of steps an agent must take in an MDP, starting from any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, to visit all state-action pairs at least once with a probability of at least 0.5.
@@
Given a state space $\mathcal{S}$, action space $\mathcal{A}$, suboptimality error $\epsilon>0$ (from Definition 2) and $10<\delta<$ 1, an algorithm $\mathcal{A}$ is called PAC-MDP (Kakade, 2003), if the number of time steps required to find a $\epsilon$-optimal policy is less than some polynomial in the relevant quantities $(|\mathcal{S}|,|\mathcal{A}|,\frac{1}{\epsilon},\frac{1}{1-\gamma},\frac{1}{\delta})$ with probability at least $1-\delta$.
For simplicity, when we say an algorithm $\mathcal{A}$ has polynomial sample complexity, we imply that $\mathcal{A}$ is PAC-MDP. The work by Liu & Brunskill (2018) establishes polynomial sample complexity for a uniformly random exploration by bounding the covering length defined in Definition 1. Using this, and considering a limited tree budget $N$, we show that $\epsilon t$-greedy is PAC-MDP. Let’s denote the search tree by $\mathcal{X}$, and the distribution over the generated options in $\mathcal{X}$ as $\mathcal{P}_{{\mathcal{X}}}$. The following Theorem provides a lower bound on option sampling in tree $\mathcal{X}$ under certain condition.
@@
[…]
5:    $s_{\min}$leftarrow root
@@
10:         if $n(\phi(s_{x^{\prime}}))$=0 then
11:             Extract option $o$ by actions $root$ to $s_{x^{\prime}}$
@@
14:         if $n(\phi$($s_{x^{\prime}})$) $<$ $n(\phi$($s_{\min})$then
15:             $s_{\min}$=$s_{x^{\prime}}$
@@
19:    Extract option $o$ by actions $root$ to $s_{\min}$
@@
[…]
@@ 3.3 Using Longest $n$-step Return
Here, $s_{T}$ is the last state in the episode. Using the longest n-step return for each transition from a successful episode, the reward is immediately propagated to all $Q$-value updates. In unsuccessful episodes, using the longest $n$-step return reduces the overestimation bias in Q-values (Thrun & Schwartz, 1993). Meng et al. (2021) empirically show that using multi-step updates can improve the performance of DDPG on robotic tasks mostly by reducing overestimation bias — they demonstrate that the larger the number of steps, the lower the estimated target Q-value and overestimation bias.
@@ 4.1 Overall Performance of ETGL-DDPG
We evaluate the performance of ETGL-DDPG compared to state-of-the-art methods. We compare with SAC (Haarnoja et al., 2018), TD3 (Fujimoto et al., 2018), DDPG, and DOIE (Lobel et al., 2022). DOIE demonstrates state-of-the-art performance in challenging sparse-reward continuous control problems by drastically improving the exploration. While both DOIE and $\epsilon t$-greedy use a similarity measure between new and observed states, DOIE applies this to compute an optimistic value function rather than solely guiding the agent to unexplored areas. The results are shown in Figure 3. In the navigation environments, ETGL-DDPG and DOIE demonstrate superior performance compared to other methods, with ETGL-DDPG achieving a success rate of 1 faster than DOIE. Notably, Wall-maze presents a more challenging task among navigation environments, where only ETGL-DDPG and DOIE are able to achieve a success rate above zero. In manipulation tasks, the press-button poses the hardest challenge as none of the methods achieve a success rate of 1. ETGL-DDPG still outperforms all other approaches, while DOIE underperforms compared to SAC, indicating its limitations in adapting to high dimensional environments.
@@ 4.2 Environment Coverage through Exploration
In $\epsilon z$-greedy, various distributions can be used to select options based on their length. Although a uniform distribution is a straightforward choice, alternative distributions such as the zeta distribution $z(n)\propto n^{-\mu}$ (where $\mu$ controls the decay rate) can also be employed. Empirical results by (Dabney et al., 2020) show that the zeta distribution slightly outperforms the uniform distribution, and moreover, produces a pattern akin to Lévy flights observed in certain ecological foraging models (Baronchelli & Radicchi, 2013). Figure 5 shows the distribution of option lengths generated by $\epsilon t$-greedy during training across all environments, revealing two main observations: (1) moderate-length options have the highest probability of being selected, and (2) probabilities decay as option length increases. While this decay pattern resembles that of the zeta distribution (where length 1 is most likely), $\epsilon t$-greedy instead favors moderate lengths. Note that Although options longer than 12 do occur, their frequency is extremely low and thus not visible in Figure 5.
@@ 5 Related Work
Other Exploration Techniques. There are also other strategies to improve exploration without using reward bonuses. Colas et al. (2018) use a policy search process to generate diverse data to improve training of DDPG. Liu et al. (2018) introduce a competition-based exploration method where two agents (A and B) compete with each other. Agent A is penalized for visiting states visited by B, while B is rewarded for visiting states discovered by A. Plappert et al. (2018) directly inject noise into the policy’s parameter space instead of the action space. Eysenbach et al. (2019) build a graph using states in the replay buffer, allowing the agent to navigate distant regions of the environment by applying Dijkstra’s algorithm. $\epsilon t$-greedy shares similarities with Eysenbach et al. (2019) in constructing a tree structure within the state space. However, while Eysenbach et al. (2019) builds a single global graph to connect the start state to the goal state, $\epsilon t$-greedy generates local trees dynamically, focusing on accessing less-explored areas. Ecoffet et al. (2021) introduce Go-Explore which separates exploration and policy optimization into two distinct phases. During the exploration phase, the algorithm systematically explores the environment by maintaining an archive of visited states along with their corresponding trajectories. In the robustification phase, these trajectories are leveraged to train a robust policy using standard reinforcement learning methods, such as PPO. However, a key limitation of Go-Explore is its reliance on resetting trajectories to return to previously visited states, a feature that is often unavailable in many real-world scenarios. Lobel et al. (2022) present Deep Optimistic Initialization for Exploration (DOIE), which improves exploration in continuous control tasks by maintaining optimism in state-action value estimates. Lobel et al. (2023) demonstrate that DOIE can estimate visit counts by averaging samples from the Rademacher distribution instead of using density models. Dey et al. (2024) present COIN, a continual optimistic initialization strategy that extends DOIE to stochastic and non-stationary environments.
@@
Experience Replay Buffer. Rather than uniformly sampling from the buffer, Prioritized Experience Replay (PER) (Schaul et al., 2015) prioritizes transitions in the buffer based on reward, recency, or TD error at the expense of $O(\log N)$ per sample, where $N$ is the buffer size. CER (Zhang & Sutton, 2017) includes the last transition from the buffer to each sampled batch with $O(1)$ complexity. Zhang et al. (2022) learn a conservative value regularizer only from the observed transitions in the replay buffer to improve the sample efficiency of DQN. Pan et al. (2022) theoretically show why PER has a better convergence rate than uniform sampling policy when minimizing mean squared error. Furthermore, Pan et al. (2022) identify two limitations of PER: outdated priorities and insufficient coverage of the state space. Kompella et al. (2023) propose Stratified Sampling from Event Tables (SSET), a method that partitions the experience replay buffer into distinct Event Tables. Each Event Table captures significant subsequences of optimal behavior, allowing for more targeted sampling during training. GDRB aligns conceptually with SSET if achieving a goal is defined as the sole event. However, SSET employs FIFO for inserting new data into its buffers, whereas GDRB uses reservoir sampling for its success-event buffer. Moreover, in SSET, the sampling ratios between buffers are fixed, while GDRB dynamically adjusts these ratios over time to favor the smaller buffer.
@@
Reward Propagation. Reward shaping (Laud, 2004; Hu et al., 2020) creates artificial intermediate rewards to facilitate reward propagation. However, designing appropriate intermediate rewards is hard and often problem-specific. Trott et al. (2019) address this issue by introducing  self-balancing reward shaping in the context of on-policy learning. Their approach defines the local optimum as the closest visited state to the goal and calculates the reward for each state based on its distance from the local optimum and the goal. By dynamically adjusting the local optimum, the agent is guided toward the optimal behavior. To extract more information from an unsuccessful episode, Andrychowicz et al. (2017) introduce imaginary goals. An imaginary goal for state $s$ is a state that is encountered later in the episode. Learning to achieve these imaginary goals helps the agent understand the structure of the environment. Memarian et al. propose Self-Supervised Online Reward Shaping (SORS), a method that alternates between ranking trajectories using sparse rewards and training a classifier to infer a dense reward function from these rankings. The inferred dense rewards are then utilized to update the agent’s policy. Devidze et al. (2024) present a reward informativeness criterion that adaptively constructs interpretable reward functions based on the agent’s current policy.

### v3

@@ 1 Introduction
Deep deterministic policy gradient (DDPG) (Lillicrap et al., 2015) is one of the representative algorithms for reinforcement learning (RL) (Sutton and Barto, 2018), alongside other prominent approaches (Haarnoja et al., 2018; Fujimoto et al., 2018; Andrychowicz et al., 2017). In recent years, DDPG has served as the backbone algorithm for introducing novel ideas in robotics and RL Barth-Maron et al. (2018b); Pan et al. (2020); Liu et al. (2023); Wang et al. (2023); Tiapkin et al. (2024). While DDPG demonstrates strong performance in continuous control tasks with dense reward signals (Duan et al., 2016; Kiran et al., 2021), its effectiveness diminishes significantly in sparse-reward settings where rewards are only observed upon reaching the goal (Matheron et al., 2019; Luck et al., 2019).
@@
In this paper, we enhance DDPG (Lillicrap et al., 2015) to address all three aforementioned problems. We choose DDPG over SAC, as SAC’s maximum entropy framework already provides advanced exploration, and combining it with our components would introduce extra complexity, making it harder to isolate and analyze our contributions. Our first contribution is $\epsilon t$-greedy, a new temporally version of $\epsilon$-greedy that utilizes a light-weight search procedure, similar to Laud (2004), to enable more directional exploration based on the agent’s previous experience data. We show that similar to $\epsilon z$-greedy, $\epsilon t$-greedy has polynomial sample complexity in related parameters of the MDP. Our second contribution is a new goal-conditioned dual replay buffer (GDRB), that uses two replay buffers along with an adaptive sampling strategy to differentiate goal-reached and goal-not-reached experience data. These two buffers differ in retention policy, size, and the transitions they store. Our third enhancement is to replace the one-step update in DDPG with the longest $n$-step return for all transitions in an episode. Figure 1 illustrates the innovations of ETGL-DDPG. In Section 4, we evaluate the performance of ETGL-DDPG through extensive experiments on 2D and 3D continuous control benchmarks. As the proposed strategies are orthogonal, we show that each of the three strategies individually improves the performance of DDPG via a thorough ablation study. Furthermore, ETGL-DDPG outperforms current state-of-the-art methods across all tested environments.
@@ 2.1 Deep Deterministic Policy Gradient (DDPG)
To ease presentation, we adopt our notation with explicit reference to the goal state for both the critic and the actor networks in DDPG. DDPG maintains an actor $\mu(s,s_{g})$ and a critic $Q(s,a,s_{g})$. The agent explores the environment through a stochastic policy $a\sim\mu(s,s_{g})+w$, where $w$ is a noise sampled from a normal distribution or an Ornstein-Uhlenbeck process (Uhlenbeck and Ornstein, 1930). To update both actor and critic, transition tuples are sampled from a replay buffer to perform a mini-batch gradient descent. The critic is updated by a loss $L$,
@@
where $y_{t}=r_{t}+\gamma Q^{\prime}(s_{t+1},\mu^{\prime}(s_{t+1},s_{g}),s_{g})$. $Q^{\prime}$ and $\mu^{\prime}$ are the target critic and actor, respectively; their weights are soft-updated to the current weights of the main critic and actor, respectively. The actor is updated by the deterministic policy gradient algorithm (Silver et al., 2014) to maximize the estimated Q-values of the critic using loss $\mathbb{E}_{s}[Q(s,\mu(s,s_{g}),s_{g})]$.
@@ 3.1 $\epsilon t$-Greedy: Exploration With Search
To justify this exploration method, we adopt the conditions outlined in Liu and Brunskill (2018) to validate the sample efficiency of $\epsilon t$-greedy. We begin by introducing the relevant terms and then present the main theorem. Detailed definitions and proofs are provided in Appendix A.1. The key idea is to define a measure that captures the concept of visiting all state-action pairs, as outlined in Definition 1.
@@
The covering length (Even-Dar and Mansour, 2004) represents the minimum number of steps an agent must take in an MDP, starting from any state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, to visit all state-action pairs at least once with a probability of at least 0.5. We define the covering length only for discrete MDPs; for continuous MDPs, we consider a discretization of the state-action space $\mathcal{S}\times\mathcal{A}$.
@@
Given a state space $\mathcal{S}$, action space $\mathcal{A}$, suboptimality error $\epsilon>0$ (from Definition 2) and $0<\delta<1$, an algorithm $\mathcal{A}$ is called PAC-MDP (Kakade, 2003), if the number of time steps required to find a $\epsilon$-optimal policy is less than some polynomial in the relevant quantities $(|\mathcal{S}|,|\mathcal{A}|,\frac{1}{\epsilon},\frac{1}{1-\gamma},\frac{1}{\delta})$ with probability at least $1-\delta$.
For simplicity, when we say an algorithm $\mathcal{A}$ has polynomial sample complexity, we imply that $\mathcal{A}$ is PAC-MDP. The work by Liu and Brunskill (2018) establishes polynomial sample complexity for a uniformly random exploration by bounding the covering length defined in Definition 1. Using this, and considering a limited tree budget $N$, we show that $\epsilon t$-greedy is PAC-MDP. Let’s denote the search tree by $\mathcal{X}$, and the distribution over the generated options in $\mathcal{X}$ as $\mathcal{P}_{{\mathcal{X}}}$. The following Theorem provides a lower bound on option sampling in tree $\mathcal{X}$ under certain condition.
@@
2:  frontier_nodes $\leftarrow\{\}$
3:  Initialize root using $s$: $\mathit{root}\leftarrow\mathit{TreeNode}(s)$
4:  frontier_nodes $\leftarrow$ frontier_nodes $\cup$ {root};
5:  $s_{\min}$ $\leftarrow$ root
@@
10:    if $n(\phi(s_{x^{\prime}}))$=1 then $\triangleright$ For a state to be in the replay buffer, it must have been visited at least once.
11:     Extract option $o$ by actions rootroot to $s_{x^{\prime}}$
@@
14:    if $n(\phi$($s_{x^{\prime}})$) << $n(\phi$($s_{\min})$) then
15:     $s_{\min}$$s_{x^{\prime}}$
@@
19:  Extract option $o$ by actions rootroot to $s_{\min}$
@@
27:  frontier_nodes $\leftarrow$ frontier_nodes $\cup$ $s_{x^{\prime}}$}
@@ 3.3 Using Longest $n$-step Return
Here, $s_{T}$ is the last state in the episode. Using the longest n-step return for each transition from a successful episode, the reward is immediately propagated to all $Q$-value updates. In unsuccessful episodes, using the longest $n$-step return reduces the overestimation bias in Q-values (Thrun and Schwartz, 1993). Meng et al. (2021) empirically show that using multi-step updates can improve the performance of DDPG on robotic tasks mostly by reducing overestimation bias — they demonstrate that the larger the number of steps, the lower the estimated target Q-value and overestimation bias.
@@ 4.1 Overall Performance of ETGL-DDPG
We evaluate the performance of ETGL-DDPG compared to state-of-the-art methods. We compare with SAC (Haarnoja et al., 2018), TD3 (Fujimoto et al., 2018), DDPG, and DOIE (Lobel et al., 2022). DOIE demonstrates state-of-the-art performance in challenging sparse-reward continuous control problems by drastically improving the exploration. While both DOIE and $\epsilon t$-greedy use a similarity measure between new and observed states, DOIE applies this to compute an optimistic value function rather than solely guiding the agent to unexplored areas. The results are shown in Figure 3. In the navigation environments, ETGL-DDPG and DOIE demonstrate superior performance compared to other methods, with ETGL-DDPG achieving a success rate of 1 faster than DOIE. Notably, Wall-maze presents a more challenging task among navigation environments, where only ETGL-DDPG and DOIE are able to achieve a success rate above zero. In manipulation tasks, the press-button poses the hardest challenge as none of the methods achieve a success rate of 1. ETGL-DDPG still outperforms all other approaches, while DOIE underperforms compared to SAC, indicating its limitations in adapting to high dimensional environments. Regarding training time, we observed that ETGL-DDPG requires approximately 1.5 times the training time of DDPG. This extended training duration can be attributed to two main components: $\epsilon t$-greedy and the longest $n$-step return. Among these, $\epsilon t$-greedy has a greater effect, as each step necessitates a local search.
@@ 4.2 Environment Coverage through Exploration
In $\epsilon z$-greedy, various distributions can be used to select options based on their length. Although a uniform distribution is a straightforward choice, alternative distributions such as the zeta distribution $z(n)\propto n^{-\mu}$ (where $\mu$ controls the decay rate) can also be employed. Empirical results by (Dabney et al., 2020) show that the zeta distribution slightly outperforms the uniform distribution, and moreover, produces a pattern akin to Lévy flights observed in certain ecological foraging models (Baronchelli and Radicchi, 2013). Figure 5 shows the distribution of option lengths generated by $\epsilon t$-greedy during training across all environments, revealing two main observations: (1) moderate-length options have the highest probability of being selected, and (2) probabilities decay as option length increases. While this decay pattern resembles that of the zeta distribution (where length 1 is most likely), $\epsilon t$-greedy instead favors moderate lengths. Note that Although options longer than 12 do occur, their frequency is extremely low and thus not visible in Figure 5.
@@ 5 Related Work
Other Exploration Techniques. There are also other strategies to improve exploration without using reward bonuses. Colas et al. (2018) use a policy search process to generate diverse data to improve training of DDPG. Liu et al. (2018) introduce a competition-based exploration method where two agents (A and B) compete with each other. Agent A is penalized for visiting states visited by B, while B is rewarded for visiting states discovered by A. Plappert et al. (2018) directly inject noise into the policy’s parameter space instead of the action space. Eysenbach et al. (2019) build a graph using states in the replay buffer, allowing the agent to navigate distant regions of the environment by applying Dijkstra’s algorithm. $\epsilon t$-greedy shares similarities with Eysenbach et al. (2019) in constructing a tree structure within the state space. However, while Eysenbach et al. (2019) builds a single global graph to connect the start state to the goal state, $\epsilon t$-greedy generates local trees dynamically, focusing on accessing less-explored areas. Ecoffet et al. (2021) introduce Go-Explore which separates exploration and policy optimization into two distinct phases. During the exploration phase, the algorithm systematically explores the environment by maintaining an archive of visited states along with their corresponding trajectories. In the robustification phase, these trajectories are leveraged to train a robust policy using standard reinforcement learning methods, such as PPO. However, a key limitation of Go-Explore is its reliance on resetting trajectories to return to previously visited states, a feature that is often unavailable in many real-world scenarios. Lobel et al. (2022) present Deep Optimistic Initialization for Exploration (DOIE), which improves exploration in continuous control tasks by maintaining optimism in state-action value estimates. Lobel et al. (2023) demonstrate that DOIE can estimate visit counts by averaging samples from the Rademacher distribution instead of using density models. Dey et al. (2024) present COIN, a continual optimistic initialization strategy that extends DOIE to stochastic and non-stationary environments. Chakraborty et al. (2023) leverage heavy-tailed action distributions to enhance exploration in continuous control tasks. Wang et al. (2022) propose a framework in which learned high-level policies select among a set of pre-designed low-level base controllers, allowing the agent to decompose complex long-horizon tasks into simpler control primitives. Dawood et al. (2023) use model predictive control (MPC), where a learned high-level policy proposes goals or waypoints, and MPC is used to generate low-level actions that guide the agent toward those goals.
@@
Experience Replay Buffer. Rather than uniformly sampling from the buffer, Prioritized Experience Replay (PER) (Schaul et al., 2015) prioritizes transitions in the buffer based on reward, recency, or TD error at the expense of $O(\log N)$ per sample, where $N$ is the buffer size. CER (Zhang and Sutton, 2017) includes the last transition from the buffer to each sampled batch with $O(1)$ complexity. Zhang et al. (2022) learn a conservative value regularizer only from the observed transitions in the replay buffer to improve the sample efficiency of DQN. Pan et al. (2022) theoretically show why PER has a better convergence rate than uniform sampling policy when minimizing mean squared error. Furthermore, Pan et al. (2022) identify two limitations of PER: outdated priorities and insufficient coverage of the state space. Kompella et al. (2023) propose Stratified Sampling from Event Tables (SSET), a method that partitions the experience replay buffer into distinct Event Tables. Each Event Table captures significant subsequences of optimal behavior, allowing for more targeted sampling during training. GDRB aligns conceptually with SSET if achieving a goal is defined as the sole event. However, SSET employs FIFO for inserting new data into its buffers, whereas GDRB uses reservoir sampling for its success-event buffer. Moreover, in SSET, the sampling ratios between buffers are fixed, while GDRB dynamically adjusts these ratios over time to favor the smaller buffer.
@@
Reward Propagation. Reward shaping (Laud, 2004; Hu et al., 2020) creates artificial intermediate rewards to facilitate reward propagation. However, designing appropriate intermediate rewards is hard and often problem-specific. Trott et al. (2019) address this issue by introducing  self-balancing reward shaping in the context of on-policy learning. Their approach defines the local optimum as the closest visited state to the goal and calculates the reward for each state based on its distance from the local optimum and the goal. By dynamically adjusting the local optimum, the agent is guided toward the optimal behavior. To extract more information from an unsuccessful episode, Andrychowicz et al. (2017) introduce imaginary goals. An imaginary goal for state $s$ is a state that is encountered later in the episode. Learning to achieve these imaginary goals helps the agent understand the structure of the environment. Memarian et al. propose Self-Supervised Online Reward Shaping (SORS), a method that alternates between ranking trajectories using sparse rewards and training a classifier to infer a dense reward function from these rankings. The inferred dense rewards are then utilized to update the agent’s policy. Devidze et al. (2024) present a reward informativeness criterion that adaptively constructs interpretable reward functions based on the agent’s current policy. Wilcox et al. (2022) propose a Monte Carlo Augmented Actor-Critic (MCA2C) framework that integrates Monte Carlo returns from suboptimal demonstration trajectories into the critic update to improve value estimation in sparse-reward settings. By combining temporal-difference targets with demonstration-based Monte Carlo returns, the method provides more informative learning signals during early training, enabling the agent to better propagate sparse rewards. Barth-Maron et al. (2018a) propose D4PG by utilizing distributional critic updates, $n$-step returns, and distributed training across multiple agents, which leads to improvements in wall-clock training time and data efficiency.